2. The table below shows the output and the total cost of a firm producing wireless

earphones. The firm charges MYR 13 per unit of output. Use this table to answer the

questions which follow. (Show your workings.)

Output

(units)

0 1,000 2,000 3,000 4,000 5,000 6,000

Total

Cost

5,000 13,000 18,000 24,000 32,000 45,000 60,000

i. Calculate the fixed cost and the variable cost when output is 3,000

units.

ii. Calculate the average variable cost when output is 5,000 units.

iii. Calculate total profit if 4,000 units are sold.

iv. Using the data in the table above, draw one graph showing the average

cost and the marginal cost of the firm, labelling them AC and MC.


1
Expert's answer
2021-11-22T10:03:36-0500

Solution:

i.). When output is 3,000 units:

Fixed cost = 5,000

Variable cost = (24,000 – 5,000) = 19,000

 

ii.). When output is 5,000 units:

Average Variable Cost (AVC) = "\frac{VC}{Q} = \frac{40,000}{5,000} = 8"

 

iii.). Profit if 4,000 units are sold:

Profit = TR – TC

Profit = (4,000 "\times"13) – 32,000 = 52,000 – 32,000 = 20,000

Profit = 20,000



 

iv.). The graph showing the average cost and the marginal cost of the firm, labeling them AC and MC are as below: